solve the system for real values of x and y
6x-y=-5
x^2+y^2=123+10y

Answer by Jk22(389)

a) 6x-y=-5
b) x^2+y^2=123+10y
a) y=6x+5
b) x^2+(y-5)^2-25=123 => x^2+(6x)^2=148=37*x^2
=> x^2=4
=> x=+2, -2
and y=17, -7 (with a)